# Robustness-Aware 3D Object Detection in Autonomous Driving: A Review and Outlook

### v1

@@ VI Conclusion
3D object detection plays a crucial role in autonomous driving perception. In recent years, this field has witnessed rapid development, yielding a plethora of research papers. Based on the diverse data forms generated by sensors, these methods are primarily categorized into three types: image-based, point cloud-based, and multimodal. The primary metrics for evaluation in these methods are high accuracy and low latency. Numerous reviews have summarized these approaches, focusing predominantly on the core principles of ‘high accuracy and low latency’ in delineating their technical trajectories. However, in the transition of autonomous driving technology from breakthroughs to practical applications, existing reviews have not prioritized safety perception as central concerns, failing to encompass the current technological pathways related to safety perception. For instance, recent multimodal fusion methods typically undergo robustness testing during the experimental phase, a facet not adequately considered in current reviews. Therefore, in this study, we reexamine 3D object detection algorithms with a central focus on the key aspects of ‘Accuracy, Latency, and Robustness.’ We reclassify previous reviews, placing particular emphasis on resegmenting from the perspective of safety perception. We aim for this work to offer new insights for future research in 3D object detection, transcending the confines of high-accuracy exploration.

### v2

@@ VII Conclusion
3D object detection plays a crucial role in autonomous driving perception. In recent years, this field has witnessed rapid development, yielding many research results. Based on the diverse data forms generated by sensors, these methods are primarily categorized into three types: image-based, point cloud-based, and multi-modal. The primary metrics for evaluation in these methods are high accuracy and low latency. Numerous reviews have summarized these approaches, focusing on the core principles of ‘high accuracy and low latency’ in delineating their technical trajectories. However, in the transition of autonomous driving technology from breakthroughs to practical applications, existing reviews have not prioritized safety perception as a central concern, failing to encompass the current technological pathways related to safety perception. For instance, recent Multi-modal fusion methods typically undergo robustness testing during the experimental phase, a facet not adequately considered in current reviews. Therefore, in this study, we re-examine 3D object detection algorithms with a central focus on the key aspects of ‘Accuracy, Latency, and Robustness’. We reclassify previous reviews, placing particular emphasis on re-segmenting from the perspective of safety perception. We aim for this work to offer new insights for future research in 3D object detection, transcending the confines of high-accuracy exploration.